### v1

Abstract
The growth problem in Gaussian elimination (GE) remains a foundational question in numerical analysis and numerical linear algebra. Wilkinson resolved the growth problem in GE with partial pivoting (GEPP) in his initial analysis from the 1960s, while he was only able to establish an upper bound for the GE with complete pivoting (GECP) growth problem. The GECP growth problem has seen a spike in recent interest, culminating in improved lower and upper bounds established by Bisain, Edelman, and Urschel in 2023, but still remains far from being fully resolved. Due to the complex dynamics governing the location of GECP pivots, analysis of GECP growth for particular input matrices often estimates the actual growth rather than computes the growth exactly. We present a class of dense random butterfly matrices on which we can present the exact GECP growth. We extend previous results that established exact growth computations for butterfly matrices when using GEPP and GE with rook pivoting (GERP) to now also include GECP for particular input matrices. Moreover, we present a new method to construct random Hadamard matrices using butterfly matrices.
keywords:
Gaussian elimination, growth factor, pivoting, butterfly matrices, numerical linear algebra
{AMS}
60B20, 15A23, 65F99
1 Introduction
Gaussian elimination (GE) remains to be one of the most used approaches to solve linear systems in modern applications. For instance, GE with partial pivoting (GEPP) is the default solver in MATLAB when using the backslash operator with a general input matrix. Additionally, GE is a staple of introductory linear algebra courses (although, based on anecdotal evidence, it may not be a favorite among all students).
GE iteratively uses elimination updates below the diagonal on an initial input matrix $A\in\mathbb{R}^{n\times n}$ to transform $A$ into an upper triangular linear system, which eventually builds the matrix factorization $A=LU$, where $L$ is a unit lower triangular matrix and $U$ is upper triangular. Each GE step transforms $A=A^{(1)}$ into the upper triangular form $U=A^{(n)}$ using $n-1$ total GE steps, where $A^{(k)}$ denotes the intermediate form of $A$ that has all zeros below the first $k-1$ entries. The lower triangular matrix $L$ is built up by successively using the pivot, i.e., the top-left entry in the untriangularized system, to scale the leading row to zero out all remaining entries below the pivot. This can then be used to solve $A{\mathbf{x}}={\mathbf{b}}$ by solving the computationally simpler triangular linear systems $L{\mathbf{y}}={\mathbf{b}}$ and ${\mathbf{U}}{\mathbf{x}}={\mathbf{y}}$ using forward and backward substitutions. General nonsingular $A$ will not have a $LU$ factorization if any leading minors are singular, i.e., if $\det(A_{ij})_{i,j=1}^{k}=0$ for some $k\leq n-1$. Moreover, numerical stability of GE when using floating-point arithmetic [^1] We will use the IEEE standard model for floating-point arithmetic. is sensitive to any elimination steps that involve division by numbers close to 0. Hence, combining GE with certain pivoting schemes is preferable even if not necessary.
GEPP remains the most popular pivoting scheme, which combines a pivot search in the leading column of the untriangularized system to then also use a row transposition so that the pivot is maximal in magnitude. This results in the GEPP factorization $PA=LU$ where $P$ is a permutation matrix, with corresponding permutation
(1) $\sigma(A)=(n\ i_{n})\cdots(2\ i_{2})(1\ i_{1}),$
where $i_{k}\geq k$ is defined by $i_{k}=\min\{\operatorname*{arg\,max}_{j\geq k}|A^{(k)}_{jk}|\}$ (corresponding to the closest maximal in magnitude entry to the original pivot entry). Moreover, also $\|L\|_{\max}=1$ (using $\|C\|_{\max}=\max_{i,j}|C_{ij}|$ for the max-norm on matrices) since then $L_{ij}=A^{(j)}_{ij}/A^{(j)}_{jj}$ satisfies $|L_{ij}|\leq 1$ for all $i>j$. We will focus in this paper in particular on GE with complete pivoting (GECP), that uses both a corresponding row and column pivot to ensure the leading pivot is maximal in the entire untriangularized system (rather than just the first column). GECP factorizations then are of the form $PAQ=LU$ where both $P,Q$ are permutation matrices, for permutations similarly in form (1). Moreover, we will assume GECP uses the default lexicographic tie-breaking pivot search strategy assuming a column-major traversal, i.e., at GECP step $k$, use the corresponding row and column swaps for the location of the first maximal magnitude entry in the minimal column and then minimal row entry. This tie-breaking strategy aligns with MATLAB’s column-major storage order, which stacks matrix columns from left to right. It then performs a GEPP pivot search on the extended column to determine the maximal element.
Using floating-point arithmetic, the numerical stability of computed solutions using GE on well-conditioned matrices relies heavily on the growth factor,
$\rho(A)=\frac{\max_{k}\|A^{(k)}\|_{\max}}{\|A\|_{\max}}.$
(The growth factor uses intermediate computed matrix factors, $\overline{A^{(k)}}$ with finite precision; we will refrain from highlighting the distinction between computed and exact intermediate factors throughout, noting also that computed and actual matrix factors align in exact arithmetic.) Using GECP, then $\max_{k}\|A^{(k)}\|_{\max}=\max_{j}|U_{jj}|$ so that $\rho(A)=\max_{j}|U_{jj}|/|U_{11}|$. For a quick demonstration of how the growth factor can be used in error analysis of computed solutions, we consider a computed solution $\hat{{\mathbf{x}}}$ in floating-point arithmetic using the GE factorization $PAQ=LU$ to the linear system $A{\mathbf{x}}={\mathbf{b}}$ where $A\in\mathbb{R}^{n\times n}$ and ${\mathbf{b}}\in\mathbb{R}^{n}$. Then, as seen in [12], the relative error then satisfies
(2) $\frac{\|{\mathbf{x}}-\hat{{\mathbf{x}}}\|_{\infty}}{\|{\mathbf{x}}\|_{\infty}}\leq 4n^{2}\epsilon_{\operatorname{machine}}\kappa_{\infty}(A)\rho(A)$
where $\kappa(A)=\|A\|_{\infty}\|A^{-1}\|_{\infty}$ is the $L^{\infty}$-induced condition number of $A$ where $\epsilon_{\operatorname{machine}}$ is the machine-epsilon that denotes the minimal positive number such that the floating-point representation $\operatorname{fl}(1+\epsilon_{\operatorname{machine}})\neq 1$ (e.g., $\epsilon_{\operatorname{machine}}=2^{-52}$ using double precision). Hence, on well-conditioned linear systems, error analysis of computed solutions can focus on analysis of the corresponding growth factors.
[…]
The difficulty in the growth problem largely comes down to the increasingly complex dynamics between entries for each iterative pivot search and subsequent elimination step. Instead of considering the growth problem across all nonsingular matrices, one can then focus on the growth problem for certain structured matrices. Hadamard matrices have their own rich history with the growth problem. $H_{n}\in\{\pm 1\}^{n\times n}$ is a Hadamard matrix if $H_{n}H_{n}^{T}=n{\mathbf{I}}_{n}$, where necessarily $n=1,2$ or $n$ is a multiple of 4 (see [3] for a larger historical overview). Hadamard matrices have a wide list of applications, including optimal design theory, coding theory, and graph theory (see, for example, [12, 16], and references therein for more background). A famous open problem with Hadamard matrices is the existence of such matrices for all multiples of 4, with currently 768 being the currently smallest such order that no known Hadamard matrices have been found yet ([5] resolved the previously lower bound of 764 in 2008). A similarly famous open problem for numerical analysis involves the growth problem for Hadamard matrices, where it is believed $\rho(H_{n})=n$. This has been established for all Sylvester Hadamard matrices (see [3]), while for general Hadamard matrices this has only been proven up to $n=16$ [15]. The growth problem for Hadamard matrices is a sub-question for the growth problem when restricted to orthogonal matrices. For instance, the orthogonal growth problem remains open in GEPP even, which was recently explored in [20].
We are interested in studying the Hadamard growth problem, as viewed in a continuous analogue version using butterfly matrices. (See Section 2 for a definition of butterfly matrices.) Butterfly matrices were introduced by D. Stott Parker as a means to remove the need of pivoting altogether using GE. For example, to solve $A{\mathbf{x}}={\mathbf{b}}$, one can use independent and identically distributed (iid) random butterfly matrices $U,V$ to instead solve the equivalent linear systems $UAV^{*}{\mathbf{y}}=U{\mathbf{b}}$ and ${\mathbf{x}}=V^{*}{\mathbf{y}}$, but now using GE with no pivoting (GENP) [18]. The primary utility in using butterfly matrices to solve this preceding transformed linear system is their recursive structure allows matrix-vector products to be performed in $O(n\log n)$ FLOPS; so these transformations can be performed without impacting the $O(n^{3})$ leading order complexity of GE itself.
In [21], the authors studied the stability properties of these butterfly transformed systems for particular initial linear systems by studying their corresponding computed growth factors when using GENP, GEPP, GE with rook pivoting (GERP, that uses iterative column and then row searches for a pivot candidate that maximizes both its column and row) and GECP. Moreover, they provide a full distributional description of the GENP, GEPP, and GERP growth factors for a particular subclass of random butterfly matrices themselves. One of our main contributions of this paper is to now extend this full distributional description to now also include GECP growth factors for a further subclass of these butterfly matrices. We also explore structural properties of GECP factorizations using different tie-breaking pivot search strategies. Additionally, we re-contextualize the discussion of butterfly matrices in terms of their growth maximizing properties for a further subclass of scaled Hadamard matrices, which we will refer to as butterfly Hadamard matrices. We then introduce a direct means to sample random butterfly Hadamard matrices.
1.1 Outline of results
The main theoretical contribution of this work is Theorem 2.3, which provides a class of butterfly matrices that do not need any row or column pivots when using GECP, i.e., they are completely pivoted. Additionally, this provides the distribution for particular random growth factors, associated to random butterfly matrices that preserve a monotonicity property for their input angle vector. This extends a previous result from [21] that provided a full distributional description of the GEPP growth factor for particular butterfly matrices. Section 2 is organized to first give sufficient background and highlight particular properties of random butterfly matrices, which is then utilized in the main technical approach given in Proposition 2.8 to maximize growth at intermediate GECP steps. Section 2.3 provides additional supporting experiments that consider GECP factorizations for butterfly matrices with different initial input orders, as well as with and without using GECP with an additional tolerance parameter for determining intermediate pivot candidates when using floating-point arithmetic. Section 3 focuses on a subclass of scaled Hadamard matrices within the butterfly matrices, called the butterfly Hadamard matrices. This class represents a rich subclass of Hadamard matrices (of order $2^{O(N\log N)}$), that necessarily maximizes the growth factor in Theorem 2.6 for a further subclass of simple scalar butterfly matrices. Moreover, we present a simple method to construct a particular butterfly Hadamard matrix for a given input random butterfly matrix.
2 Growth factors of random butterfly matrices
Butterfly matrices are a family of recursive orthogonal transformations that were introduced by D. Stott Parker in 1995 as a means of accelerating common computations in linear algebra [18]. We will use a particular family of butterfly matrices built up using rotation matrices, studied in more detail previously in [21, 27]. We will now review particular definitions of these butterfly matrices, along certain GE matrix factorization forms and growth factor results for particular butterfly matrices.
2.1 Random butterfly matrices
An order $N=2^{n}$ butterfly matrix $B$ takes the recursive form
(3) $B=\begin{bmatrix}C&S\\ -S&C\end{bmatrix}\begin{bmatrix}A_{1}&{\mathbf{0}}\\ {\mathbf{0}}&A_{2}\end{bmatrix}=\begin{bmatrix}CA_{1}&SA_{2}\\ -SA_{1}&CA_{2}\end{bmatrix},$
using symmetric and commuting matrices, $C,S$, that satisfy the matrix Pythagorean identity $C^{2}+S^{2}={\mathbf{I}}_{N/2}$ and $A_{1},A_{2}$ are order $N/2$ butterfly matrices, where we define $B=1$ if $N=1$. If $A_{1}=A_{2}$ at each recursive step, then the resulting butterfly matrices are called simple butterfly matrices. By construction, necessarily $B\in\operatorname{SO}(N)$ and also $(C,S)$ have corresponding eigenvalue pairs $(\cos\theta,\sin\theta)$ for some angle $\theta$. The primary utility for butterfly matrices is due to the fast matrix-vector multiplication afforded by this recursive structure, which uses $O(Nn)$ FLOPS assuming $C,S$ have $O(N)$ matrix-vector multiplication complexity; this can be ensured by then using diagonal $C,S$ at each recursive step.
Let $\operatorname{B}(N)$ and $\operatorname{B}_{s}(N)$ denote the order $N$ scalar nonsimple and simple butterfly matrices, formed using scalar $(C_{k},S_{k})=(\cos\theta_{k},\sin\theta_{k}){\mathbf{I}}_{2^{k-1}}$ at each recursive step $k$ for some angle $\theta_{k}$. Similarly, let $\operatorname{B}^{(D)}(N)$ and $\operatorname{B}_{s}^{(D)}(N)$ denote the order $N$ diagonal nonsimple and simple butterfly matrices, formed using diagonal $(C_{k},S_{k})=\bigoplus_{j=1}^{2^{k-1}}(\cos\theta_{k}^{(j)},\sin\theta_{k}^{(j)})$ for each recursive step $k$.
Recall the Kronecker product of $A\in\mathbb{R}^{n_{1}\times m_{1}},B\in\mathbb{R}^{n_{2}\times m_{2}}$ is the matrix $A\otimes B\in\mathbb{R}^{n_{1}n_{2}\times m_{1}m_{2}}$ of the form
$A\otimes B=\begin{bmatrix}A_{11}B&\cdots&A_{1m_{1}}B\\ \vdots&\ddots&\vdots\\ A_{n_{1}1}B&\cdots&A_{n_{1}m_{1}}B\end{bmatrix}.$
We will also use the notation $A\oplus B\in\mathbb{R}^{(n_{1}+n_{2})\times(m_{1}+m_{2})}$ to denote the block diagonal matrix with diagonal blocks $A,B$. Using these notations, $B_{1}\in\operatorname{B}(N)$ and $B_{2}\in\operatorname{B}_{s}(N)$ have (3) take the forms
$B_{1}=(R_{\theta}\otimes{\mathbf{I}}_{N/2})(A_{1}\oplus A_{2}),\qquad B_{2}=\bigotimes_{j=1}^{n}R_{\theta_{j}},\qquad\mbox{with}\qquad R_{\theta}=\begin{bmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{bmatrix}$
denoting the standard (clockwise) rotation matrices. Further recall the Kronecker product satisfies the mixed-product property, where
$(A\otimes B)(C\otimes D)=(AC)\otimes(BD)$
when matrix dimensions are compatible. Moreover, both $\otimes$ and $\oplus$ preserve certain matrix structures and operators, such as permutation matrices, triangular forms, inverses, and transposes. Additionally, there exist perfect shuffle permutations (see [4]) $P,Q$ such that $P(A\otimes B)Q=B\otimes A$; if $A,B$ are both square, then $Q=P^{T}$. Hence, for $Q_{n}$ the perfect shuffle matrices such that $Q_{n}(A\otimes{\mathbf{I}}_{N/2})Q_{n}^{T}={\mathbf{I}}_{N/2}\otimes A$ for $A\in\mathbb{R}^{2\times 2}$, then the diagonal butterfly matrices $B_{1}\in\operatorname{B}^{(D)}(N)$ and $B_{2}\in\operatorname{B}^{(D)}_{s}(N)$ have (3) take the form
(4) $B_{1}=Q_{n}\left(\bigoplus_{j=1}^{N/2}R_{\theta_{j}}\right)Q_{n}^{T}\cdot(A_{1}\oplus A_{2})$
for $A_{1},A_{2}\in B^{(D)}(N/2)$, where additionally $A_{1}\oplus A_{2}={\mathbf{I}}_{2}\otimes A$ in the simple case $A_{1}=A_{2}=A\in B^{(D)}_{s}(N/2)$. A straightforward check confirms the number $M$ of input angles needed to generate a butterfly matrix from $\operatorname{B}_{s}(N),\operatorname{B}(N),\operatorname{B}_{s}^{(D)}(N),\operatorname{B}^{(D)}(N)$ is, respectively, $n$, $N-1$, $N-1$, and $nN/2$. We will write $B(\boldsymbol{\theta})$ with $\boldsymbol{\theta}\in[0,2\pi)^{M}$ then for each respective type of butterfly matrix.
Random butterfly matrices are formed using random input angles. Let $\Sigma$ be a collection of generating pairs $(C_{k},S_{k})$ satisfying the above butterfly properties with random input angles. Then $\operatorname{B}(N,\Sigma)$ and $\operatorname{B}_{s}(N,\Sigma)$ will denote the ensembles of random butterfly matrices and random simple butterfly matrices formed by constructing a butterfly matrix by independently sampling $(C,S)$ from $\Sigma$ at each recursive step. Let
$\Sigma_{S}$ =$=\{(\cos\theta^{(k)},\sin\theta^{(k)}){\mathbf{I}}_{2^{k-1}}:\theta^{(k)}\mbox{ iid }\operatorname{Uniform}([0,2\pi)),k\geq 1\}\quad\mbox{and}$
$\Sigma_{D}$ =$=\{\bigoplus_{j=1}^{2^{k-1}}(\cos\theta_{j}^{(k)},\sin\theta_{j}^{(k)}):\theta^{(k)}_{j}\mbox{ iid }\operatorname{Uniform}([0,2\pi)),k\geq 1\}.$
Note from the mixed-product property, $\operatorname{B}_{s}(N)=\bigotimes^{n}\operatorname{SO}(2)$ is an abelian subgroup of $\operatorname{SO}(N)$, and hence has a Haar probability measure. Using uniform input angles then leads to induced uniform measures on the push-forward transformations, so that:
Proposition 2.1 ([27]).
$\operatorname{B}_{s}(N,\Sigma_{S})\sim\operatorname{Haar}(\operatorname{B}_{s}(N))$.
We will then refer to $\operatorname{B}_{s}(N,\Sigma_{S})$ as the Haar-butterfly matrices, as this explicit construction provides a method to directly sample from this distribution.
A natural question using butterfly transformations on linear systems is to determine how far away is the transformed system from the input system. Since butterfly matrices are orthogonal, then multiplication then is equivalent to taking a one-step orthogonal walk with initial location at the input matrix; so multiplying by more butterfly matrices then corresponds to a longer orthogonal walk.
We can measure how far a step is in this orthogonal walk by just viewing how much does perturbing the input angles change the overall matrix. This can be answered explicitly by computing bounds for $\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}$, which show the map $\boldsymbol{\theta}\mapsto B(\boldsymbol{\theta})$ is Lipschitz continuous.
Proposition 2.2.
The maps $\boldsymbol{\theta}\mapsto B(\boldsymbol{\theta})$ for $B(\boldsymbol{\theta})$ in $\operatorname{B}_{s}(N),\operatorname{B}(N),\operatorname{B}_{s}^{d}(N)$ and $\operatorname{B}^{d}(N)$ are each Lipschitz continuous with respect to the $2$-norm to Frobenius normed spaces, with respective Lipschitz constants of $\sqrt{N}$, $\sqrt{2(N-1)}$, $\sqrt{2(N-1)}$ and $\sqrt{2n}$.
In particular, this shows the butterfly models comprise connected manifolds of dimension $M$ for each respective type of butterfly matrices. A proof of Proposition 2.2 is found in Appendix A.
2.2 GECP growth factor of butterfly matrices
The growth factor is an important statistic that controls the precision when using GE to solve linear systems, as seen in (2). In particular, for well-conditioned linear systems, then error analysis on GE can focus on analysis of the growth factor itself. The study of growth factors on random linear systems has its own rich history.
The previous focus on worst-case behavior to gauge algorithm applications led earlier researchers to focus more on GECP when using GE, due to its sub-exponential worst-case upper bound, and hence the ongoing interest in the GECP growth problem. However, despite potentially exponential growth, typical GEPP growth remained much smaller in practice. Understanding why this is the case has remained a continued question of interest in numerical analysis.
A good starting point on the study of random growth factors is the average-case analysis of Schreiber and Trefethen in 1990 [26]. Their analysis shifted the focus of the growth problem from worst-case behavior to instead look at the average behavior on different random matrix ensembles. Using different pivoting schemes, they presented empirical estimates of the first moment on iid random Gaussian ensembles of size up to $10^{4}$, showing GEPP typically maintained polynomial growth factors of order $n^{2/3}$ while GECP had growth of order $n^{1/2}$. Understanding the behavior of growth factors on fixed linear systems using additive Gaussian perturbations falls under the heading of smoothed analysis (see, e.g., [24] for a general overview of using smoothed analysis to analyze algorithms). A large step forward in understanding the overview behavior of GE was accomplished by Sankar, Speilman, and Teng who provided a full smoothed analysis of GENP in 2006 [23]. However, the increasing complexity of pivoting prevented their methods from being transferable to a full smoothed analysis for GEPP. For instance, only partial smoothed analysis results of GEPP were carried out by Sankar in his Ph.D. thesis [22]. Huang and Tikhomirov gave the most significant contribution to improving our understanding of GEPP in their improved average-case analysis using iid Gaussian matrices, where they provide high probability polynomial bounds on the growth factors [13]. Again, a full smoothed analysis of GEPP remains out of reach still, as their methods only establish improved growth factor bounds on additive Gaussian perturbations of the all zeros matrix (compared to Sankar’s previous work). Other recent work has explored dynamics between GEPP and GECP growth factors, as well as studying growth behavior using multiplicative random orthogonal perturbations [20].
While previous results using random growth factors focused on empirical moment estimates and bounds, the authors in [21] take advantage of the structural properties of Haar-butterfly matrices to then provide full distributional descriptions for certain growth factors, including GENP, GEPP, and GERP. A similar approach was taken in [8] for improved lower bounds on the GERP growth problem, where Kronecker products were further utilized as they preserve certain maximal growth properties. (However, this does not extend to GECP.)
We recall the statement of the theorem (cf. [21, Theorem 2.3]) that resolves the GENP, GEPP, and GERP growth problem for Haar-butterfly matrices:
Theorem 2.3 ([21]).
Let $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$ and $X\sim\operatorname{Cauchy}(1)$. Then $\rho(B)\sim\prod_{j=1}^{n}(1+Y_{j}^{2})$ where $Y_{j}$ are iid, with $Y_{j}\sim|X|$ when using GENP and $Y_{j}\sim(|X|\mid|X|\leq 1)$ when using GEPP or GERP. Moreover, using any pivoting scheme then $\rho(B)\geq\rho^{\operatorname{GEPP}}(B)$.
The proof of Theorem 2.3 utilized $\tan\theta\sim X$ and $\max(|\tan\theta|,|\cos\theta|)\sim(|X|\mid|X|\leq 1)$ for $X\sim\operatorname{Cauchy}(1)$ when $\theta\sim\operatorname{Uniform}([0,2\pi))$, in addition to certain maximizing pivoting properties among the intermediate GE forms of $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$, which can be derived from the following lemma:
Lemma 2.4 ([21]).
Suppose $A\in\mathbb{R}^{N/2\times N/2}$ has an $LU$ factorization using GENP. Let
$B=\begin{bmatrix}\cos\theta A&\sin\theta A\\ -\sin\theta A&\cos\theta A\end{bmatrix}=R_{\theta}\otimes A$
for $\cos\theta\neq 0$. Then $B$ has an $LU$ factorization using GENP. Moreover, if $k\leq N/2$, then
(5) $B^{(k)}=\begin{bmatrix}\cos\theta A^{(k)}&\sin\theta A^{(k)}\\ -\sin\theta\begin{bmatrix}{\mathbf{0}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{I}}_{N/2-k+1}\end{bmatrix}A^{(k)}&\sec\theta\left(A-\sin^{2}\theta\begin{bmatrix}{\mathbf{0}}&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{I}}_{N/2-k+1}\end{bmatrix}A^{(k)}\right)\end{bmatrix}.$
If $k=N/2+j$ for $j\geq 1$, then
(6) $B^{(k)}=\begin{bmatrix}\cos\theta A^{(N/2)}&\sin\theta A^{(N/2)}\\ {\mathbf{0}}&\sec\theta A^{(j)}\end{bmatrix}.$
In particular, Lemma 2.4 then yields
(7) $\rho(B(\boldsymbol{\theta}))=\prod_{j=1}^{n}\max(|\sec\theta_{j}|,|\csc\theta_{j}|)^{2}=\prod_{j=1}^{n}(1+\max(|\tan\theta_{j}|,|\cot\theta_{j}|)^{2})$
for $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ if using GEPP (or GERP). It follows then this product is invariant under any permutation of the indices for the input angles:
Corollary 2.5.
Let $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ and $\sigma\in S_{n}$. If $\tilde{\theta}_{j}=\theta_{\sigma(j)}$, then $\rho(B(\boldsymbol{\theta}))=\rho(B(\tilde{\boldsymbol{\theta}}))$.
(a) $P$
(b) $L+U$
(c) $Q$
Figure 1: Sparsity patterns for (computed, within the given tolerance $\operatorname{tol}=10^{4}\cdot\epsilon_{\operatorname{machine}}$) matrix factors (a) $P$, (b) $L+U$, and (c) $Q$ from then GECP factorization $PBQ=LU$ for $B=B(\tilde{\boldsymbol{\theta}})\in\operatorname{B}_{s}(N)$ for $N=2^{10}$ and $\boldsymbol{\theta}\sim\operatorname{Uniform}([0,2\pi)^{10})$.
Our goal is now to extend Theorem 2.3 to include a subclass of butterfly matrices when using also GECP. In particular, we will introduce a class of matrices that do not need any pivoting when using GECP, i.e., they are completely pivoted:
Theorem 2.6.
Let $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$. If $\boldsymbol{\theta}\in[0,2\pi)^{n}$ such that
(8) $|\tan\theta_{j+1}|\leq|\tan\theta_{j}|\leq 1$
for each $j$, then the GENP, GEPP, GERP, and GECP factorizations of $B({\boldsymbol{\theta}})$ all align, i.e, $B(\boldsymbol{\theta})$ is completely pivoted. In particular, then the growth factor takes the form (7). Moreover, if $\boldsymbol{\theta}\sim\operatorname{Uniform}([0,2\pi)^{n})$ and $\sigma\in S_{n}$ is such that $\tilde{\theta}_{j}=\theta_{\sigma(j)}$ where
(9) $\max(|\tan\tilde{\theta}_{j+1}|,|\cot\tilde{\theta}_{j+1}|)\leq\max(|\tan\tilde{\theta}_{j}|,|\cot\tilde{\theta}_{j}|),$
then the GEPP, GERP, and GECP factorizations of $B(\tilde{\boldsymbol{\theta}})$ all align and $\rho(B(\tilde{\boldsymbol{\theta}}))$ has distribution determined by Theorem 2.3.
Figure 1 shows plots of the sparsity patterns for the (computed) factors $P$, $L+U$, and $Q$ where $PBQ=LU$ is the GECP factorization for $B=B(\tilde{\boldsymbol{\theta}})\in\operatorname{B}_{s}(N)$ with $N=2^{10}$, using $\boldsymbol{\theta}\sim\operatorname{Uniform}([0,2\pi)^{n})$. Note the $L$ and $U$ factors then have sparsity patterns determined by the Kronecker factor forms of $L$ and $U$, that then have nonzero entry locations aligning for $L$ and $U^{T}$ in line with the Sierpiński triangle, determined by $\bigotimes^{n}\begin{bmatrix}1&0\\ 1&1\end{bmatrix}$; this follows directly from the mixed-product property and the $LU$ factorization
(10) $R_{\theta}=\begin{bmatrix}1&0\\ -\tan\theta&1\end{bmatrix}\begin{bmatrix}\cos\theta&\sin\theta\\ 0&\sec\theta\end{bmatrix}:=L_{\theta}U_{\theta}.$
The $\max(|\tan\theta|,|\cot\theta|)$ terms in Theorem 2.6 arise in terms of the corresponding Kronecker matrix factors from the overall GEPP factorization along with the mixed-product property (noting further $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ has GEPP factorization determined by the Kronecker product of the individual GEPP factorizations for each Kronecker factor). Furthermore, note $|\tan\theta|\geq 1$ is equivalent to no GEPP pivot movement being needed with $R_{\theta}$ since $|\cos\theta|\geq|\sin\theta|$, while $\tan\theta|<1$ (or $|\cot\theta|>1$) aligns with the case when a GEPP pivot movement is needed with $R_{\theta}$. Moreover, the map of $Q$ in Figure 1 aligns with the identity matrix, which shows no column pivots were needed, while $P$ aligns with the associated simple butterfly permutation that would arise using GEPP on $B$. Other properties of simple butterfly permutations, including those that determine the number of GEPP pivot movements needed with $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$, are studied more extensively in [19]. Future work will additionally study other properties of the sparsity patterns of $P,Q$ in the context of permuton theory (cf. [2]).
(a) $P$
(b) $L+U$
(c) $Q$
Figure 2: Sparsity patterns for the computed GECP factors of $B$ from Figure 1, now using GECP without an added tolerance parameter to identify potential candidates during each intermediate pivot search.
Remark 2.7.
Figure 1 used an additional tolerance level of $\operatorname{tol}=10^{4}\cdot\epsilon_{\operatorname{machine}}$ to determine the nonzero entries from $L+U$. This tolerance was also used for a custom MATLAB GECP algorithm to ensure the lexicographic column-major ordering was enforced, after first determining all possible pivot candidates within this tolerance level of the computed overall maximal entry in the remaining untriangularized block. Figure 2 shows the sparsity patterns of the same $B$ from Figure 1, but now using a GECP algorithm without the added tolerance to determine the full list of potential pivot candidates. This in particular highlights how the floating-point arithmetic errors accumulate in such a way so that the computed “maximal” entry then manifests as a “random” candidate among what would be potential pivot candidates. Of note, although the overall sparsity pattern was not preserved without the tolerance parameter, the actual sparsity (i.e., proportion of zero entries, which still is using the tolerance to distinguish what a “zero” entry is) does remain intact. Additionally, the $L$ and $U^{T}$ factors as well appear to maintain a symmetric sparsity pattern (as is confirmed at least with our sample since $L$ and $U^{T}$ had entries with magnitude larger than the tolerance in the exact same locations).
To prove Theorem 2.6, we will prove the following proposition that that encapsulates the maximizing growth properties of the particular matrices considered in Theorem 2.6. This is an extension and update to a similar proposition used in [21] to establish GEPP (and GERP) butterfly maximizing growth properties.
Proposition 2.8.
Let $B=B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ such that $|\tan\theta_{i+1}|\leq|\tan\theta_{i}|\leq 1$ for all $i$, and let $B=LU$ be the $LU$ factorization of $B$ using GENP. Let $\eta,\varepsilon\in\mathbb{R}$ such that $|\varepsilon|\leq|\eta-\varepsilon|$. Then for all $k$
(11) $\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}_{N-k+1}\end{bmatrix}(\eta B-\varepsilon B^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}_{N-k+1}\end{bmatrix}\right\|_{\max}\leq|\eta-\varepsilon||U_{kk}|.$
In particular, then $B=LU$ is also the $LU$ factorization of $B$ using GECP, i.e., Theorem 2.6 holds.
Proof 2.9.
First, note the last implication follows since then
$|B^{(k)}_{kk}|\leq\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}_{N-k+1}\end{bmatrix}B^{(k)}\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}_{N-k+1}\end{bmatrix}\right\|_{\max}\leq|U_{kk}|=|B^{(k)}_{kk}|$
for all $k$ using (11) with $\eta=0$ and $\varepsilon=1$, so that no row or column swaps would be needed at any intermediate GECP step. This determines the first part of Theorem 2.6, while the latter half follows then also from Theorem 2.3 and Corollary 2.5.
To prove (11), we will use induction on $n=\log_{2}N$. Note first the result is immediate for $k=1$ since then
$\|{\mathbf{I}}_{N}(\eta B-\varepsilon B^{(1)}){\mathbf{I}}_{N}\|_{\max}=|\eta-\varepsilon|\|B\|_{\max}=|\eta-\varepsilon||U_{11}|.$
So it suffices to assume $k\geq 2$. If $n=1$, then using (10) with $B^{(2)}=U=U_{\theta}$, we see for $k=2$ then
$\left\|\begin{bmatrix}0&1\end{bmatrix}(\eta B-\varepsilon B^{(2)})\begin{bmatrix}0\\ 1\end{bmatrix}\right\|_{\max}=|\eta\cos\theta-\varepsilon\sec\theta|=|\eta\cos^{2}\theta-\varepsilon||U_{22}|\leq|\eta-\varepsilon||U_{22}|$
using Lemma 2.10 stated below, with $\alpha=\beta=1$.
Lemma 2.10.
Let $\alpha,\beta,\theta,\eta,\varepsilon\in\mathbb{R}$ such that $|\varepsilon|\leq|\eta-\varepsilon|$. Then
$|\eta\alpha\cos^{2}\theta-\varepsilon\beta|\leq\cos^{2}\theta|\eta\alpha-\varepsilon\beta|+\sin^{2}\theta|\eta-\varepsilon||\beta|$
Proof 2.11.
Since $\eta\alpha\cos^{2}\theta-\varepsilon\beta=\cos^{2}\theta(\eta\alpha-\varepsilon\beta)-\sin^{2}\theta\varepsilon\beta,$ this follows directly by the triangle inequality and $|\varepsilon|\leq|\eta-\varepsilon|$.
[…]
(12) $U_{kk}=\left\{\begin{array}[]{ll}\cos\theta U_{kk}^{\prime}&\mbox{if $k\leq N/2$}\\ \sec\theta U_{jj}^{\prime}&\mbox{if $k=N/2+j$ for $j\geq 1$.}\end{array}\right.$
For $21<k\leq N/$ 2, then for ${\mathbf{I}}={\mathbf{I}}_{N/2-k+1}$ when not indicated otherwise, we have
$\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}_{N-k+1}\end{bmatrix}(\eta B-\varepsilon B^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}_{N-k+1}\end{bmatrix}$
$\hskip 12.0pt=\begin{bmatrix}\cos\theta\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta A-\varepsilon A^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}&\sin\theta\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta A-\varepsilon A^{(k)})\\ -\sin\theta\left(\eta A-\varepsilon\begin{bmatrix}{\mathbf{0}}\\ &{\mathbf{I}}\end{bmatrix}A^{(k)}\right)\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}&-\sec\theta\left((\varepsilon-\eta\cos^{2}\theta)A-\varepsilon\sin^{2}\theta\begin{bmatrix}{\mathbf{0}}\\ &{\mathbf{I}}\end{bmatrix}A^{(k)}\right)\end{bmatrix}$
using Lemma 2.4. Let $\eta^{\prime}=\varepsilon-\eta\cos^{2}\theta$ and $\varepsilon^{\prime}=\varepsilon\sin^{2}\theta$ where we note
$|\varepsilon^{\prime}|=\sin^{2}\theta|\varepsilon|\leq\cos^{2}\theta|\eta-\varepsilon|=|\eta^{\prime}-\varepsilon^{\prime}|$
since $\sin^{2}\theta\leq\cos^{2}\theta$ and $|\varepsilon|\leq|\eta-\varepsilon|$. By the inductive hypothesis, we have
$\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta^{\prime}A-\varepsilon^{\prime}A^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}$ $\leq|\eta^{\prime}-\varepsilon^{\prime}||U^{\prime}_{kk}|=|\cos\theta||\eta-\varepsilon||U_{kk}|\quad\mbox{and}$
$\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta A-\varepsilon A^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}$ $\leq|\eta-\varepsilon||U_{kk}^{\prime}|=|\sec\theta||\eta-\varepsilon||U_{kk}|$
using (12). Also, since $|\cos\theta_{i+1}|\geq|\cos\theta_{i}|$ for all $i$, then
$|\sec\theta|\|A\|_{\max}=|\sec\theta_{n}||U_{11}^{\prime}|=\frac{\prod_{j=1}^{n-1}|\cos\theta_{j}|}{|\cos\theta_{n}|}\leq\frac{\prod_{j\in\mathcal{J}_{k}}|\cos\theta_{j}|}{\prod_{j\in[n]\backslash\mathcal{J}_{k}}|\cos\theta_{j}|}=|U_{kk}|$
for some $\varnothing\neq\mathcal{J}_{k}\subset[n]$ when $k>1$ by Lemma 2.4. Next, we see
$|\eta\sin\theta|\|A\|_{\max}\leq|\eta-\varepsilon||2\sin\theta|\|A\|_{\max}\leq|\eta-\varepsilon||\sec\theta|\|A\|_{\max}\leq|\eta-\varepsilon||U_{kk}|$
using $|\eta|\leq|\eta-\varepsilon|+|\varepsilon|\leq 2|\eta-\varepsilon|$ and $|2\sin\theta|\leq|\sec\theta|$ (since $|2\sin\theta\cos\theta|=|\sin(2\theta)|\leq 1)$. It follows
$\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}_{N-k+1}\end{bmatrix}(\eta B-\varepsilon B^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}_{N-k+1}\end{bmatrix}\right\|_{\max}$
$\hskip 12.0pt=\max\left(\begin{array}[]{c}|\eta\sin\theta|\|A_{N/2-k+1:,:k-1}\|_{\max},\\ |\eta\sin\theta|\|A_{:k-1,N/2-k+1:}\|_{\max},\\ |\cos\theta|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta A-\varepsilon A^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max},\\ |\sec\theta||\varepsilon-\eta\cos^{2}\theta|\|A\|_{\max},\\ |\sec\theta|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}\left((\varepsilon-\eta\cos^{2}\theta)A-\varepsilon\sin^{2}\theta A^{(k)}\right)\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}\end{array}\right)$
$\hskip 12.0pt\leq|\eta-\varepsilon||U_{kk}|$
using again Lemma 2.10 so that $|\varepsilon-\eta\cos^{2}\theta|\leq|\eta-\varepsilon|$.
For $k=N/2+j$ and $j\geq 1$ so that $N-k+1=N/2-j+1$, writing now ${\mathbf{I}}={\mathbf{I}}_{N-k+1}={\mathbf{I}}_{N/2-j+1}$, we have
$\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta B-\varepsilon B^{(k)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}$
$\hskip 12.0pt=|\sec\theta|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta\cos^{2}\theta A-\varepsilon A^{(j)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}$
$\hskip 12.0pt=|\sec\theta|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}\left|\eta\cos^{2}\theta A-\varepsilon A^{(j)}\right|\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}$
$\hskip 12.0pt\leq|\sec\theta|\left(\cos^{2}\theta\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}\left|\eta A-\varepsilon A^{(j)}\right|\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}+\sin^{2}\theta|\eta-\varepsilon|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}|A^{(j)}|\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}\right)$
$\hskip 12.0pt=|\sec\theta|\left(\cos^{2}\theta\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}(\eta A-\varepsilon A^{(j)})\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}+\sin^{2}\theta|\eta-\varepsilon|\left\|\begin{bmatrix}{\mathbf{0}}&{\mathbf{I}}\end{bmatrix}A^{(j)}\begin{bmatrix}{\mathbf{0}}\\ {\mathbf{I}}\end{bmatrix}\right\|_{\max}\right)$
$\hskip 12.0pt\leq|\sec\theta||\eta-\varepsilon||U^{\prime}_{jj}|=|\eta-\varepsilon||U_{kk}|$
using Lemma 2.4 for the first equality, Lemma 2.10 for the first inequality (applied componentwise with $\alpha=A_{i^{\prime}j^{\prime}}$ and $\beta=A^{(j)}_{i^{\prime}j^{\prime}}$), the inductive hypothesis for the last inequality (with $\eta=0$ and $\varepsilon=1$ for the second term), (12) for the last equality, and the fact $\|A\|_{\max}=\||A|\|_{\max}$ for the remaining steps.
Theorem 2.6 extends a classic result from Tornheim from 1970 that produced particular Sylvester Hadamard matrices that were complete pivoted:
Proposition 2.12 ([25]).
If $A$ is completely pivoted and $H=\begin{bmatrix}1&1\\ -1&1\end{bmatrix}$, then $A\otimes H$ is completely pivoted. In particular, then $\bigotimes^{n}H$ is completely pivoted.
There is a natural relationship between butterfly matrices and particular Sylvester Hadamard matrices, which will be further outlined in Section 3. Of note, butterfly matrices can be interpreted as continuous generalizations of these Sylvester Hadamard matrices, on which these Hadamard matrices then achieve certain maximizing properties for particular statistics, including the growth factors. Proposition 2.12 only guarantees $B({\boldsymbol{\theta}})\otimes H$ remains completely pivoted if $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ is completely pivoted. Since $H=\sqrt{2}\cdot B([\frac{\pi}{4},\frac{\pi}{4}]^{T})$, and matrices remain completely pivoted under multiplication by scalar matrices and diagonal sign matrices (e.g., if $A\in\mathbb{R}^{n\times n}$ is completely pivoted, then $A\cdot(\bigoplus^{n}\pm 1)$ is also completely pivoted), then Theorem 2.6 further guarantees $B(\boldsymbol{\theta})\otimes R_{\varphi}$ is completely pivoted if $|\tan\varphi|\geq\max_{j}|\tan\theta_{j}|$ rather than requiring only $|\tan\varphi|=1$. Moreover, as noted in [3], Proposition 2.12 is not symmetric: if $A$ is completely pivoted, then $H\otimes A$ is not necessarily completely pivoted; this can be seen by considering $A=\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$. Theorem 2.6 has the additional flexibility that $R_{\varphi}\otimes B(\boldsymbol{\theta})$ is completely pivoted if $|\tan\varphi|\leq\min_{j}|\tan\theta_{j}|$, as well as forming additionally completely pivoted butterfly matrices by embedding Kronecker rotation matrix factors that preserve the ordering from Theorem 2.6.
2.3 Experiments for computed permutation factors
A natural question from Theorem 2.6 is whether the additional structure of $\operatorname{B}_{s}(N)$ lends itself so that GECP factorizations are attainable using any ordering of input angles. Unfortunately, this happens to only be the case for small $n$.
[…]
Table 1 shows the output using computed GECP factorizations (using the added tolerance parameter for an expanded pivot candidate search) for $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$, counting only those where the reordered input angles produce a permutation factor other than those formed using the perfect shuffles $\tilde{Q}_{n}$ and the GEPP butterfly permutation. (See [6, 19] for the particular perfect shuffle and butterfly permutation forms then possible.) The set of experiments used multiple samples first of $\boldsymbol{\theta}\sim\operatorname{Uniform}([0,2\pi)^{n})$, transforming each angle with $\hat{\theta}=\theta$ if $|\tan\theta|\leq 1$ (i.e., no GEPP pivot movement is needed on $R_{\theta}$) or $\hat{\theta}=\frac{\pi}{2}-\theta$ if $|\tan\theta|>1$ (i.e., a GEPP pivot movement is needed on $R_{\theta}$), then reordering the transformed angles $\hat{\boldsymbol{\theta}}$ into $\tilde{\boldsymbol{\theta}}$ so that $B(\tilde{\boldsymbol{\theta}})$ is completely pivoted (i.e., satisfies the hypothesis of Theorem 2.6). Then the GECP factorization (using the added tolerance of $10^{3}\cdot\epsilon_{\operatorname{machine}}$ to determine an initial list of potential pivot candidates) is computed for each of the $|S_{n}|=n!$ rearrangements of the angles of $\tilde{\boldsymbol{\theta}}$. Table 1 then records the count per $\sigma\in S_{n}$ rearrangement $\sigma(\boldsymbol{\theta})$ such that $\sigma(\boldsymbol{\theta})_{j}=\tilde{\boldsymbol{\theta}}_{\sigma(j)}$ of $\tilde{\boldsymbol{\theta}}$ only if either the computed $P,Q$ factors differ from the corresponding perfect shuffle $\tilde{Q}_{n}$ to return $PB(\sigma(\tilde{\boldsymbol{\theta}}))Q=B(\tilde{\boldsymbol{\theta}})$. The experiment is then repeated (between 2 to 5 times) for differing initial $\boldsymbol{\theta}\sim\operatorname{Uniform}([0,2\pi)^{n})$, recording any instance where a computed GECP factorization does not align with the corresponding perfect shuffle rearrangement.
$n$ $|S_{n}|$$\#\{PBQ\neq(\tilde{Q}_{n})B(\tilde{Q}_{n}^{T})\}$
1 1 - -
2 2 - -
3 6 - -
4 24 2 8.3%
5 120 36 30.0%
6 720 381 52.9%
Table 1: Computed number of GECP factorizations $PBQ=LU$ for $B\in\operatorname{B}_{s}(N)$ where $P,Q$ take a form other than that expected from the corresponding perfect shuffle and GEPP factorization.
It happens that for $n=1,2,3$, then the only permutation factors encountered are precisely those of the form $P=\tilde{Q}_{n}$ and $Q=\tilde{Q}_{n}^{T}$. However, this is no longer the case starting at $n=4$. As $n$ grows, then the proportion of arrangements of input angles that align with the needed perfect shuffle permutation shrink with $n$, where $n=6$ is the smallest such size that leads to at least half of the rearrangements differing from the perfect shuffle.
(a) $P$
(b) $L+U$
(c) $Q$
(d) $P$
(e) $L+U$
(f) $Q$
Figure 3: Sparsity patterns for the computed GECP factors of $B\in\operatorname{B}_{s}(2^{10})$ not satisfying Theorem 2.6, with (a)-(c) corresponding to GECP with an added tolerance parameter during the pivot candidate compilation, and (d)-(f) without the added tolerance setting.
For example, if $\boldsymbol{\theta}\in[0,2\pi)^{4}$ such that $|\tan\theta_{4}|<|\tan\theta_{3}|<|\tan\theta_{2}|<|\tan\theta_{1}|\leq 1$, then the computed GECP factorization using the reordered input angles $[\theta_{4},\theta_{3},\theta_{1},\theta_{2}]^{T}$ takes the form $P_{\sigma_{1}}B([\theta_{4},\theta_{3},\theta_{1},\theta_{2}]^{T})P_{\sigma_{2}}^{T}$, where
[…]
compared to the perfect shuffle permutation $\rho$ with $\tilde{Q}_{n}=P_{\rho}$, where
[…]
In particular, considering the given forms (1) forms for each permutation, then the first GECP step where the computed permutation differed from the perfect shuffle was at step 5. (So the first 4 GECP steps align for both strategies.) Two pivot candidates at GECP step 5 include the entries at locations $(6,6)$ and $(9,9)$. Using the column-major lexicographic GECP tie-breaking strategy, then the $(6,6)$ is deemed the new pivot and so each computed permutation uses the row and column transpositions (5 6)56(5\ 6)( 5 6 ). In order to have recovered the necessary perfect shuffle, the GECP tie-breaking strategy would have needed to choose instead the entry at $(9,9)$ and hence the transposition (5 9)59(5\ 9)( 5 9 ).
Figure 3 shows both the sparsity patterns for the GECP factors of $B$ not in the particular ordered form found in Theorem 2.6, first showing the GECP with the added tolerance parameter (to widen the potential pivot candidates before implementing the column-major lexicographic tie-breaking strategy), which shows the Kronecker product is broken for this particular input $B\in\operatorname{B}_{s}(2^{10})$. This also includes the sparsity patterns for the GECP factors without using the added tolerance parameter for the pivot search, that essentially returns a “random” pivot among the potential pivot candidates due to rounding. Again, an interesting observation is the sparsity and symmetry of the final $L+U$ terms is preserved still, both with and without the added tolerance.
(a) $\rho(B)$
(b) $\rho_{\infty}(B)$
Figure 4: Overlaid histograms of computed growth factors (a) $\rho(B)$ and (b) $\rho_{\infty}(B)$ where $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$, using GEPP and both GECP with and without an added tolerance parameter to enforce the column-major lexicographic tie-breaking strategy, using $10^{4}$ trials with $N=2^{8}$ and a logarithmic scaling.
Although Theorem 2.6 yields the explicit random growth factors for butterfly matrices whose input angles preserve a particular monotonicity property, we expect the random growth factors using GECP to not actually need to rely on the input ordering. Theorem 2.6 used this ordering to then determine every explicit intermediate GECP form of the butterfly linear system, including explicitly the final matrix factorization $PBQ=LU$, that could then determine how the local growth properties are presented at each step. Theorem 2.3 guaranteed the growth factor using GE with any pivoting scheme is minimized using GEPP. This further established that GERP also attained this minimal value using any ordering of the input angles, as the GERP and GEPP factorizations then aligned perfectly (still assuming a column-major GERP pivot search). Empirically, we expect this to also be the case with GECP. Figure 4 shows the results of running $10^{4}$ trials, uniformly sampling $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$ with $N=2^{8}$ (a sufficient size chosen for efficient sampling). For each sampled matrix, we computed the GEPP growth factors, along with each GECP growth factors using first the added tolerance parameter (with $\operatorname{tol}=10^{3}\cdot\epsilon_{\operatorname{machine}}$) to ensure the column-major lexicographic tie-breaking strategy was enforced, as well as the GECP growth factors without this added tolerance parameter. In all $10^{4}$ trials, the computed growth factors matched exactly within $7.1054\cdot 10^{-14}$. Figure 4(a) shows the overlaid final histograms of the computed growth factors $\rho(B)$ on a logarithmic scaling. For comparison, an alternative growth factor
$\rho_{\infty}(A)=\frac{\|L\|_{\infty}\|U\|_{\infty}}{\|A\|_{\infty}}$
was additionally computed for each GE factorization. This was used in particular on the experiments in [21], which presented an analogous full distributional description for $\rho_{\infty}(B)$ for $B\sim\operatorname{B}_{s}(N,\Sigma_{S})$ when using GEPP. Unlike using the max-norm $\|\cdot\|_{\max}$ (as in $\rho(B)$ and seen in Figure 4(a)), the induced $L^{\infty}$ matrix norm $\|\cdot\|_{\infty}$ was more sensitive to the corresponding row and column pivots encountered in GECP as seen in the non-overlaying histograms in Figure 4(b). Of note, the added randomness of the location of the pivot at each intermediate step in the GECP without tolerance led to a smaller overall computed $L^{\infty}$-growth factor.
3 Butterfly Hadamard matrices
As evidenced in (7), the growth factors for butterfly matrices are maximized when $\max(|\tan\theta_{j}|,|\cot\theta_{j}|)=|\tan\theta_{j}|=1$ for each $j$, in which case $\rho(B)=2^{n}=N$. Hence, $\rho(B(\boldsymbol{\theta}))\in\operatorname{B}_{s}(N)$ is maximal whenever $\boldsymbol{\theta}\in\{\pm\frac{\pi}{2}\pm\frac{\pi}{4}\}=\{\frac{\pi}{4},\frac{3\pi}{4},\frac{5\pi}{4},\frac{7\pi}{4}\}$. In this case, since then $\sqrt{2}\cdot R_{\theta}=H\cdot(\pm 1\oplus\pm 1)$; it follows $H(\boldsymbol{\theta})=\sqrt{N}\cdot B(\boldsymbol{\theta})$ is a scaled Hadamard matrix. A similar construction works for any scalar or diagonal butterfly matrix formed only using input angles ±$\pm\frac{\pi}{2}\pm\frac{\pi}{4}$, such that $H(\boldsymbol{\theta})=\sqrt{N}\cdot B(\boldsymbol{\theta})$ is again a Hadamard matrix; this will be formally established in Proposition 3.4. For convention, we will refer to such Hadamard matrices as butterfly Hadamard matrices, and we will let $\operatorname{H}_{s}(N),\operatorname{H}(N),\operatorname{H}^{(D)}_{s}(N),\operatorname{H}^{(D)}(N)$ then correspond to the butterfly Hadamard matrices formed respectively by requiring $B(\boldsymbol{\theta})$ belong to $\operatorname{B}_{s}(N),\operatorname{B}(N),\operatorname{B}_{s}^{(D)}(N),\operatorname{B}^{(D)}(N)$.
The most famous ongoing conjecture with Hadamard matrices corresponds to the existence of such a matrix of all orders divisible by 4. The smallest such order for which no Hadamard matrix has been found (yet) is $768$ [5]. Hadamard matrices have their own rich history in terms of the growth problem. The $n^{O(\log n)}$ GECP growth bound Wilkinson provided was believed to be far from optimal. For nearly 30 years, a popular belief was that the growth was actually bounded by $n$, and that this could be achieved only by Hadamard matrices (see [8] for further historical background). This was disproved by Edelman in [7] when he confirmed a $13\times 13$ matrix with growth of about 13.0205, upgrading a previous empirical counterexample using floating-point arithmetic one year earlier by Gould in [11] to a true counterexample in exact arithmetic. However, the previous conjectured growth bound of $n$ still remains intact for Hadamard matrices (and orthogonal matrices; see [20]).
The Hadamard growth problem further subdivides the target to consider additionally the growth problem on different equivalence classes of Hadamard matrices: two Hadamard matrices $H_{1},H_{2}$ are equivalent if there exists two signed permutation matrices $P_{1},P_{2}$ such that $P_{1}H_{1}=H_{2}P_{2}$. There is only one Hadamard equivalence class for orders up to 12, while there are 5 distinct classes of order 16, 2 for order 20, 60 for order 24, 487 for 28, and then millions for the next few orders (e.g., 13,710,027 for 32; see [14]). By considering all Hadamard equivalence classes, the Hadamard growth problem has only been resolved up to $n=16$ [15]. The Hadamard growth problem is additionally resolved (using Proposition 2.12) for Sylvester equivalent Hadamard matrices of all orders. A straightforward check shows that $H(\boldsymbol{\theta})\in\operatorname{H}_{s}(N)$ is equivalent to a Sylvester Hadamard matrix, so that then Theorem 2.6 provides an additional resolution path for the Sylvester Hadamard growth problem. In this context, as well as with Proposition 2.2, then the simple scalar butterfly growth problem can be viewed as a continuous generalization of the Sylvester Hadamard growth problem.
Another open problem of interest with Hadamard matrices involves the number of such possible matrices of a given size $m$. By considering only the $m!$ permutations of the rows, then a lower bound on the count of Hadamard matrices is $2^{\Omega(m\log m)}$ while an upper bound of $2^{\binom{m+1}{2}}$ is also easy to attain (see [9] for an overview of this conjecture, along with the current improved upper bound of order $2^{(1-c)m^{2}/2}$ for a fixed constant $c$; it is further conjectured the lower bound is the correct asymptotic scaling). The butterfly Hadamard matrices produce a rich set of Hadamard matrices, which then can be used as a simple way to produce random Hadamard matrices when using random input angles from ±$\pm\frac{\pi}{2}\pm\frac{\pi}{4}$ (e.g., consider a uniform angle from this set), as is seen here:
Proposition 3.1.
For $N=2^{n}$, then $|\operatorname{H}_{s}(N)|=2^{n+1}=2N$, $|\operatorname{H}(N)|=2^{3\cdot 2^{n-1}-1}=2^{\frac{3}{2}N-1}$, $|\operatorname{H}_{s}^{(D)}(N)|=2^{2^{n}-n+1}=2^{N-n+1}$, and $|\operatorname{H}^{(D)}(N)|=2^{2^{n-1}n+1}=2^{Nn/2+1}$. In particular, $|H^{(D)}(N)|=2^{O(N\log N)}$ matches the conjectured asymtpotic bound on the number of Hadamard matrices.
Proof 3.2.
If $\alpha_{n}$ is the respective size for each set of butterfly Hadamard matrices of order $2^{n}$, then each satisfy $\alpha_{1}=4$ along with the respective recurrence relations
$\alpha_{n+1}=\left\{\begin{array}[]{ll}\alpha_{1}\cdot\alpha_{n}/2=2\alpha_{n}&\mbox{for $\operatorname{H}_{s}(N)$,}\\ \alpha_{1}\cdot\alpha_{n}^{2}/2=2\alpha_{n}^{2}&\mbox{for $\operatorname{H}(N)$,}\\ \alpha_{1}^{N/2}\cdot\alpha_{n}/2=2^{2^{n}-1}\alpha_{n}&\mbox{for $\operatorname{H}_{s}^{d}(N)$, and}\\ \alpha_{1}^{N/2}\cdot\alpha_{n}^{2}/2=2^{2^{n}-1}\alpha_{n}^{2}&\mbox{for $\operatorname{H}^{d}(N)$,}\end{array}\right.$
where (3) can be used to first determine the number of different inputs for each of the left and right factors, which then overcount by 2 as the number of ways that the identity matrix can be factored out (with $-{\mathbf{I}}$ terms from both the left and right). A direct check then verifies each of the above counts satisfy each respective recurrence. (Equivalently, one can solve the recurrences using instead $\beta_{n}=\log_{2}\alpha_{n}$.)
Remark 3.3.
A different way to equivalently count $|\operatorname{H}_{s}(N)|$ is to first consider the $4^{n}$ possible input vectors with $\theta_{j}=\pm\frac{\pi}{2}\pm\frac{\pi}{4}$. This overcounts by all inputs that have scalar factors that combine to produce the identity matrix. This occurs with an even number of the input Kronecker factors can have a $-{\mathbf{I}}_{2}$ that factors out, which occurs $\sum_{k\operatorname{even}}\binom{n}{k}=2^{n-1}$ times. Hence, the total count is then $4^{n}/2^{n-1}=2^{n+1}=2N$.
Popular constructions of Hadamard matrices include the Sylvester and Walsh-Hadamard constructions that work for for $N=2^{n}$ or $N=mk$ where a Hadamard matrix is known to exist of order $m$ and $k$. Alternative constructions of Hadamard matrices include the Paley construction, which uses tools from finite fields, and the Williamson construction, which uses the sum of four squares [17, 31]. Applications of these only work for very particular orders and do not cover all potential multiple of 4 orders.
The butterfly Hadamard matrices provide an additional method of constructing certain Walsh-Hadamard matrices. As previously mentioned, $\operatorname{H}_{s}(N)$ yield particular Sylvester equivalent Hadamard matrices. This gives one way to generate a Sylvester Hadamard matrix using the butterfly matrices, with $2N$ possible butterfly Hadamard matrices of order $N$. In light of Proposition 2.2, the particular butterfly Hadamard matrices can also be interpreted as being Hadamard representations for butterfly matrices whose input vectors have $(\cos\theta_{j},\sin\theta_{j})$ with particular sign combinations in {(±$\{(\pm 1,\pm 1)\}^{M}$, with $M=n,N-1,N-1,Nn/2$ for respectively scalar simple, scalar nonsimple, diagonal simple, and diagonal nonsimple butterfly matrices (assuming $|\cos\theta_{j}\sin\theta_{j}|\neq 0$, so that $\theta_{j}$ is not an integer multiple of $\frac{\pi}{2}$). Hence, the butterfly Hadamard matrices can be viewed as representing the fixed sector (±$(\pm 1,\pm 1)^{M}$ on which the input angle vector maps to.
Rather than fixing given angles, one can generate a Hadamard matrix using (almost) any butterfly matrix combined with the $\operatorname{sgn}$ function, which when applied to a matrix acts componentwise so that $(\operatorname{sgn}(A))_{ij}=\operatorname{sgn}(A_{ij})=\frac{A_{ij}}{|A_{ij}|}$, where we define $\operatorname{sgn}(0):=0$. If $A\in(\mathbb{R}\backslash\{0\})^{n\times m}$ then $\operatorname{sgn}(A)\in\{\pm 1\}^{n\times m}$. In particular, note if $\theta_{j}\not\in\frac{\pi}{2}\mathbb{Z}$ for all $j$, then $\operatorname{sgn}(B(\boldsymbol{\theta}))\in\{\pm 1\}^{N\times N}$ for any scalar or diagonal butterfly matrix.
Note $\operatorname{sgn}(DA)=\operatorname{sgn}(D)\operatorname{sgn}(A)$ if $D$ is a diagonal matrix. It follows then
(13) $\operatorname{sgn}(A\otimes B)=\operatorname{sgn}(A)\otimes\operatorname{sgn}(B),\quad\operatorname{sgn}\left((A\otimes{\mathbf{I}}_{N/2})(B\oplus C)\right)=(\operatorname{sgn}(A)\otimes{\mathbf{I}}_{N/2})(\operatorname{sgn}(B)\oplus\operatorname{sgn}(C))$
for $A\in\mathbb{R}^{2\times 2}$ and $B,C\in\mathbb{R}^{N/2}$. In particular, for $C,S$ diagonal, then
(14) $\operatorname{sgn}\left(\begin{bmatrix}CA_{1}&SA_{2}\\ -SA_{1}&CA_{2}\end{bmatrix}\right)=\begin{bmatrix}\operatorname{sgn}(C)\operatorname{sgn}(A_{1})&\operatorname{sgn}(S)\operatorname{sgn}(A_{2})\\ -\operatorname{sgn}(S)\operatorname{sgn}(A_{1})&\operatorname{sgn}(C)\operatorname{sgn}(A_{2})\end{bmatrix}.$
For $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$, from (13) we have
(15) $\operatorname{sgn}(B(\boldsymbol{\theta}))=\sqrt{N}\cdot B(\hat{\boldsymbol{\theta}})\quad\mbox{where}\quad\hat{\theta}_{j}=\frac{\pi}{4}\left(2\left\lfloor\frac{2\theta_{j}}{\pi}\right\rfloor+1\right),$
where then $\hat{\theta}$ sends $\theta$ to the representative of ±$\pm\frac{\pi}{2}\pm\frac{\pi}{4}$ based on which quadrant $(\cos\theta,\sin\theta)$ is located. For example, $\hat{1}=\frac{\pi}{4}$ and $\hat{4}=\frac{5\pi}{4}$. Additionally, from (14) we have $\operatorname{sgn}(B)$ is a Walsh-Hadamard matrix for also $B$ a nonsimple scalar butterfly matrix, simple diagonal butterfly matrix and nonsimple diagonal butterfly matrix assuming $\theta_{j}\not\in\frac{\pi}{2}\mathbb{Z}$ for any $j$. The nonsimple scalar butterfly matrix case follows immediately from (14). The diagonal case is less direct:
Proposition 3.4.
If $B(\boldsymbol{\theta})$ is a scalar or diagonal order $N$ butterfly matrix with $\theta_{j}\not\in\frac{\pi}{2}\mathbb{Z}$ for all $j$, then $\operatorname{sgn}(B(\boldsymbol{\theta}))=\sqrt{N}\cdot B(\hat{\boldsymbol{\theta}})$ is a Hadamard matrix.
Proof 3.5.
The equality in $\operatorname{sgn}(B(\boldsymbol{\theta}))=\sqrt{N}\cdot B(\hat{\boldsymbol{\theta}})$ follows from Proposition 2.2 and the fact $\operatorname{sgn}$ is constant on sectors of $\mathbb{R}^{N\times N}$. To establish $H=\operatorname{sgn}(B)$ is a Hadamard matrix for $B$ a butterfly matrix, it suffices to consider only $B=B(\boldsymbol{\theta})\in\operatorname{B}^{(D)}(N)$ (since each other butterfly ensemble is a subset of this). Since $\theta_{j}\not\in\frac{\pi}{2}\mathbb{Z}$ for all $j$, then $H\in\{\pm 1\}^{N\times N}$. To see $HH^{T}=N{\mathbf{I}}_{N}$, we will use induction on $n$.
First, we will consider $(C,S)(\boldsymbol{\theta})=\bigoplus_{j=1}^{N/2}(\cos\theta_{j},\sin\theta_{j})$. Since $C$ and $S$ are nonsingular diagonal matrices, then $\operatorname{sgn}(C)^{2}=\operatorname{sgn}(S)^{2}={\mathbf{I}}_{N/2}$ and $\operatorname{sgn}(C)\operatorname{sgn}(S)=\operatorname{sgn}(S)\operatorname{sgn}(C)$. It follows
(16) $\begin{bmatrix}\operatorname{sgn}(C)&\operatorname{sgn}(S)\\ -\operatorname{sgn}(S)&\operatorname{sgn}(C)\end{bmatrix}\begin{bmatrix}\operatorname{sgn}(C)&\operatorname{sgn}(S)\\ -\operatorname{sgn}(S)&\operatorname{sgn}(C)\end{bmatrix}^{T}=2{\mathbf{I}}_{N}.$
This establishes the base case for $n=2$. Now assume the result holds for $n-1$. From (14), we have
(17) $H=\begin{bmatrix}\operatorname{sgn}(C)&\operatorname{sgn}(S)\\ -\operatorname{sgn}(S)&\operatorname{sgn}(C)\end{bmatrix}\begin{bmatrix}\operatorname{sgn}(A_{1})&{\mathbf{0}}\\ {\mathbf{0}}&\operatorname{sgn}(A_{2})\end{bmatrix}$
where $A_{1},A_{2}\in\operatorname{B}^{(D)}(N/2)$. By the inductive hypothesis, $\operatorname{sgn}(A_{i})\operatorname{sgn}(A_{i})^{T}=\frac{N}{2}{\mathbf{I}}_{N/2}$ so that
(18) $(\operatorname{sgn}(A_{1})\oplus\operatorname{sgn}(A_{2}))(\operatorname{sgn}(A_{1})\oplus\operatorname{sgn}(A_{2}))^{T}=\frac{N}{2}{\mathbf{I}}_{N}.$
Combining Eqs. 16 and 18 yields $HH^{T}=N{\mathbf{I}}_{N}$.
Remark 3.6.
The subsampled randomized Hadamard transformation (SRHT) is popular in data compression, dimensionality reduction, and random algorithms, which is often implemented with a fixed fast Walsh Hadamard matrix and a random diagonal sign matrix (see [28]). This in particular makes uses of the fast $O(N\log N)$ matrix-vector multiplication property, while the randomness is presented in the random diagonal sign matrix. Future work can explore potentially added utility of random butterfly Hadamard transformations, that maintain this quick matrix-vector multiplication but with additional randomness injected that may better serve certain structured models.
Remark 3.7.
Similar constructions using butterfly matrices initiated with $\operatorname{O}(2)$ instead of $\operatorname{SO}(2)$ generate additional Hadamard matrices, but these are still of order $2^{\mathcal{O}(N)}$.
Appendix A Connectedness of butterfly matrices
First, recall $\|\cdot\|_{F}$ is invariant under unitary transformations. Using also the triangle inequality (as seen in [10]), we have for $U_{i},V_{i}\in\operatorname{U}(N)$ then
(19) $\|U_{1}V_{1}-U_{2}V_{2}\|_{F}\leq\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F}.$
Moreover, directly by the definition of $\|\cdot\|_{F}$,
(20) $\|A\oplus B\|_{F}^{2}=\|A\|_{F}^{2}+\|B\|_{F}^{2}$
and
(21) $\|A\otimes B\|_{F}=\|A\|_{F}\|B\|_{F}.$
Next, we will explore the continuity of the butterfly factors.
Recall for $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$, then using the mixed-product property of the Kronecker product we have
(22) $B(\boldsymbol{\theta})=U_{n}\Lambda_{\boldsymbol{\theta}}U_{n}^{*}$
for unitary $U_{n}=\bigotimes^{n}U$ and $\Lambda_{\boldsymbol{\theta}}=\bigotimes_{j=1}^{n}\Lambda_{\theta_{n-j+1}}$, where
(23) $U=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ i&-i\end{bmatrix}\quad\mbox{and}\quad\Lambda_{\theta}=\begin{bmatrix}e^{i\theta}&\\ &e^{-i\theta}\end{bmatrix}.$
Since $\Lambda_{\theta+\varepsilon}=\Lambda_{\theta}\Lambda_{\varepsilon}$, then $B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})=B(\boldsymbol{\theta})B(\boldsymbol{\varepsilon})$. (In fact, it easily follows $\operatorname{B}_{s}(N)$ is an abelian subgroup of $\operatorname{SO}(N)$.) Also, recall there exist perfect shuffle permutations $P_{1},P_{2}$ such that
(24) $B\otimes A=P_{1}(A\otimes B)P_{2}.$
If $A,B$ are both square, then $P_{1}=P_{2}^{T}$. Next, recall the elementary bound
(25) $1-\cos x\leq\frac{x^{2}}{2}$
for all $x\in\mathbb{R}$, which can be expanded to yield
(26) $1-\prod_{j=1}^{n}\cos x_{j}\leq\frac{1}{2}\|{\mathbf{x}}\|_{2}^{2}$
for any ${\mathbf{x}}\in\mathbb{R}^{n}$.222(25) follows by minimizing the auxiliary function $f(x)=\frac{x^{2}}{2}+\cos x$ and (26) follows through induction along with the fact $1-xy=1-x+x(1-y)$.
This leads to the result for simple scalar butterfly matrices:
Lemma A.1.
Let $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$ and $\boldsymbol{\varepsilon}\in\mathbb{R}^{n}$. Then
(27) $\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}\leq\sqrt{N}\|\boldsymbol{\varepsilon}\|_{2}.$
Proof A.2.
Note first
(28) $B({\boldsymbol{\varepsilon}})+B({\boldsymbol{\varepsilon}})^{*}=2\prod_{j=1}^{n}\cos\varepsilon_{j}{\mathbf{I}}_{N}.$
Now to establish (27), we see
$\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}^{2}$ $=\|B(\boldsymbol{\theta})({\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon}))\|_{F}^{2}=\|{\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon})\|_{F}^{2}$
$=\operatorname{Tr}(({\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon}))({\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon}))^{*})=\operatorname{Tr}(2{\mathbf{I}}_{N}-(B(\boldsymbol{\varepsilon})+B(\boldsymbol{\varepsilon})^{*}))$
$=\operatorname{Tr}\left(2\left(1-\prod_{j=1}^{n}\cos\varepsilon_{j}\right){\mathbf{I}}_{N}\right)=2N\left(1-\prod_{j=1}^{n}\cos\varepsilon_{j}\right)$
$\leq N\|\boldsymbol{\varepsilon}\|_{2}^{2}.$
using (26) and (28) for the last two lines.
Note if $\boldsymbol{\theta}=\theta{\mathbf{e}}_{j}$ for some $j$, then $B(\boldsymbol{\theta})={\mathbf{I}}_{N2^{-j}}\otimes B(\theta)\otimes{\mathbf{I}}_{2^{j-1}}$. Hence, Lemma A.1 yields:
(29) $\|{\mathbf{I}}_{N}-{\mathbf{I}}_{N2^{-j}}\otimes B(\varepsilon)\otimes{\mathbf{I}}_{2^{j-1}}\|_{F}^{2}\leq N\varepsilon^{2}$
for each $j=1$ to $n$.
To establish similar results for $\operatorname{B}(N)$, $\operatorname{B}^{(D)}_{s}(N)$ and $\operatorname{B}^{(D)}(N)$, we can use the following multiplicative decompositions based on (3) for
(30) $\operatorname{B}(N)$ $=\prod_{j=1}^{n}\mathcal{D}(N,N2^{-j}),$
(31) $\operatorname{B}_{s}^{(D)}(N)$ $=\prod_{j=1}^{n}\mathcal{D}^{(D)}_{s}(N,N2^{-j})\quad\mbox{and}$
(32) $\operatorname{B}^{(D)}(N)$ $=\prod_{j=1}^{n}\mathcal{D}^{(D)}(N,N2^{-j})$
where the corresponding butterfly block factors, with block diagonal matrix models with $k$ blocks, are of the forms
(33) $\mathcal{D}(N,k)$ $=\bigoplus^{k}\operatorname{SO}(2)\otimes{\mathbf{I}}_{N/2k}$
(34) $\mathcal{D}^{(D)}_{s}(N,k)$ $=\mathcal{D}^{(D)}_{s}(N/k,1)\otimes{\mathbf{I}}_{k}=P_{k,N}\mathcal{D}(N,N/2k)P_{k,N}^{T},\quad\mbox{and}$
(35) $\mathcal{D}^{(D)}(N,k)$ $=\bigoplus^{k}\mathcal{D}^{(D)}_{s}(N/k,1)$
for $P_{k,N}$ the perfect shuffle permutation matrix such that $A\otimes{\mathbf{I}}_{k}=P_{k,N}({\mathbf{I}}_{k}\otimes A)P_{k,N}^{T}$ for any $A\in\mathbb{R}^{N/k\times N/k}$. In particular, $\mathcal{D}^{(D)}_{s}(N,1)$ is the general rotation matrix formed using the diagonal matrices from (4). Note $\operatorname{B}_{s}(N)$ can similarly be decomposed into a product of block butterfly factors, where each factor is of the form
(36) $\mathcal{D}_{s}(N,k)={\mathbf{I}}_{k}\otimes\operatorname{SO}(2)\otimes{\mathbf{I}}_{N/2k},$
but it is easier to deal directly with the Kronecker product factors themselves in the prior arguments.
Now we can establish a similar result for each of the block butterfly factors:
Lemma A.3.
(I) Let $B(\boldsymbol{\theta})\in\mathcal{D}(N,k)$ and $\boldsymbol{\varepsilon}\in\mathbb{R}^{k}$. Then
(37) $\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}^{2}\leq{\frac{N}{k}}\|\boldsymbol{\varepsilon}\|_{2}^{2}.$
(II) Let $B(\boldsymbol{\theta})\in\mathcal{D}^{(D)}_{s}(N,k)$ and $\boldsymbol{\varepsilon}\in\mathbb{R}^{N/2k}$. Then
$\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}^{2}\leq{2k}\|\boldsymbol{\varepsilon}\|_{2}^{2}.$
(III) Let $B(\boldsymbol{\theta})\in\mathcal{D}^{(D)}(N,k)$ and $\boldsymbol{\varepsilon}\in\mathbb{R}^{N/2}$. Then
(38) $\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}^{2}\leq 2\|\boldsymbol{\varepsilon}\|_{2}^{2}.$
Proof A.4.
Note although $\operatorname{B}(N)$, $\operatorname{B}_{s}^{(D)}(N)$ and $\operatorname{B}^{(D)}(N)$ are not themselves groups (they are not closed under multiplication), each of the butterfly block models are abelian groups (since $\operatorname{SO}(2)$ is abelian). In particular, $B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})=B(\boldsymbol{\theta})B(\boldsymbol{\varepsilon})$ for each case. It follows
(39) $\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}=\|B(\boldsymbol{\theta})({\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon}))\|_{F}=\|{\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon})\|_{F}.$
For (I), it suffices to show
(40) $\|{\mathbf{I}}_{N}-B(\boldsymbol{\varepsilon})\|_{F}^{2}$ $=\|\bigoplus_{j=1}^{k}({\mathbf{I}}_{N/k}-B(\varepsilon_{j})\otimes{\mathbf{I}}_{N/2k})\|_{F}^{2}$ $=\sum_{j=1}^{k}\|{\mathbf{I}}_{N/k}-B(\varepsilon_{j})\otimes{\mathbf{I}}_{N/2k}\|_{F}^{2}\leq\frac{N}{k}\|\boldsymbol{\varepsilon}\|_{2}^{2}$
using (20) and Eq. 29. (II) follows from (I) along with (34) and the unitary invariance of $\|\cdot\|_{F}$. (III) follows from (II) along with (35).
Now we can establish the corresponding Lipschitz constants for the maps $\boldsymbol{\theta}\mapsto B(\boldsymbol{\theta})$ as stated in Proposition 2.2.
Proof A.5 (Proof of Proposition 2.2).
The result for $\operatorname{B}_{s}(N)$ follows directly from Lemma A.1. Next, let $B(\boldsymbol{\theta})\in\operatorname{B}(N)$ for $\boldsymbol{\theta}\in\mathbb{R}^{N-1}$. Let $\boldsymbol{\theta}_{j}\in\mathbb{R}^{N2^{-j}}$ such that $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{n})$ and
$B(\boldsymbol{\theta})=B_{n}(\boldsymbol{\theta}_{n})\cdots B_{1}(\boldsymbol{\theta}_{1})=\prod_{j=1}^{n}B_{n-j+1}(\boldsymbol{\theta}_{n-j+1})$
for $B_{j}(\boldsymbol{\theta}_{j})\in\mathcal{D}(N,N2^{-j})$. Similarly decompose $\boldsymbol{\varepsilon}$ so that
$\|B(\boldsymbol{\theta})-B(\boldsymbol{\theta}+\boldsymbol{\varepsilon})\|_{F}$ $=\left\|\prod_{j=1}^{n}B_{n-j+1}(\boldsymbol{\theta}_{n-j+1})-\prod_{j=1}^{n}B_{n-j+1}(\boldsymbol{\theta}_{n-j+1}+\boldsymbol{\varepsilon}_{n-j+1})\right\|_{F}$
$\leq\sum_{j=1}^{n}\|B_{j}(\boldsymbol{\theta}_{j})-B_{j}(\boldsymbol{\theta}_{j}+\boldsymbol{\varepsilon}_{j})\|_{F}$
(41) $\leq\sum_{j=1}^{n}2^{j/2}\|\boldsymbol{\varepsilon}_{j}\|_{2}$
$\leq\left(\sum_{j=1}^{n}2^{j}\right)^{1/2}\left(\sum_{j=1}^{n}\|\boldsymbol{\varepsilon}_{j}\|_{2}^{2}\right)^{1/2}$
$=\sqrt{2(N-1)}\|\boldsymbol{\varepsilon}\|_{2}.$
using (19), (I) from Lemma A.3, and the Cauchy-Schwartz inequality, respectively, for the above inequalities. The result for $\operatorname{B}_{s}^{(D)}(N)$ follows similarly, where (41) now has $\sqrt{N2^{1-j}}$ in place of $2^{j/2}$ using (II) from Lemma A.3. The result for $\operatorname{B}^{(D)}(N)$ follows also similarly, where now in (41) we have $\sqrt{2}$ in place of $2^{j/2}$ using (III) from Lemma A.3.
Recall the Givens rotations are defined
$G(\theta,i,j)=P_{(2\ j)(1\ i)}(B(\theta)\oplus{\mathbf{I}}_{n-2})P_{(2\ j)(1\ i)}^{T}$
and can be used to generate $\operatorname{SO}(n)$ using $\frac{n(n-1)}{2}$ Givens rotations (and hence to sample $\operatorname{Haar}(\operatorname{SO}(n))$ using uniform angles). Note this exact argument can be used to verify the Lipshitz continuity of this mapping
$\boldsymbol{\theta}\mapsto A(\boldsymbol{\theta})=\prod_{i=1}^{n-1}\prod_{j=i+1}^{n}G(\theta_{\alpha_{ij}},i,j),$
for $\alpha_{ij}=\binom{n-1}{2}-\binom{n-i}{2}+j-i$ that gives an indexing of the numbers of 1 to $\binom{n-1}{2}=\frac{n(n-1)}{2}$ following the lexicographic ordering of $i,j$ in the above product. For example, if $i=1$, then $k_{1j}=j-1$ returns 1 to $n-1$, then for $i=2$, then $k_{1j}=n-1+j-1$ then returns $n$ to $(n-1)+(n-2)=2n-3$, and so forth. In fact, a direct computation establishes (as seen in [20, Lemma 4.1]):
Proposition A.6 ([20]).
The map $\boldsymbol{\theta}\mapsto A(\boldsymbol{\theta})\in\operatorname{SO}(N)$ for $\boldsymbol{\theta}\in[0,2\pi)^{m}$ with $m=\frac{N(N-1)}{2}=\dim(\operatorname{SO}(N))$ is Lipschitz continuous with constant $\sqrt{N(N-1)}$.
Remark A.7.
Note (19) can also be used to derive the Lipschitz-continuity of $\boldsymbol{\theta}\mapsto B(\boldsymbol{\theta})$ for $B(\boldsymbol{\theta})\in\operatorname{B}_{s}(N)$. However, the corresponding result would yield a Lipschitz constant of $\sqrt{Nn}$ (since $\sqrt{N}$ would be used in place of $2^{j/2}$ in (41)), which is weaker than the $\sqrt{N}$ Lipschitz constant found directly in Lemma A.1. One could then note consistently using the above argument, an in particular using (19) with respect to the decomposition as a product of block butterfly factors, yields the Lipshitz constants that are precisely the square roots of twice the inverse ranking of the number of parameters (i.e., the manifold dimension) needed to generate $\operatorname{B}_{s}(N)$, $\operatorname{B}(N)$, $\operatorname{B}_{s}^{(D)}(N)$ and $\operatorname{B}^{(D)}(N)$, i.e., respectively, $n$, $N-1$, $N-1$ and $\frac{1}{2}Nn$.